For a dipole, q=  2 × 10-6 C and d = 4 cm. The external electric field if maximum torque experienced by this dipole is 10-2 Nm, is

Grade:12

1 Answers

Khimraj
5 years ago
maximum torque experienced by this dipole is given by qdE
so 10-2 = 2*10-6*4*10-2*E
So E = 1.25*105N/C.
Hope it clears. If you like anser then please approve it.